OpenStudy (anonymous):

solve this logarithmic and express the solution in exact form log2 (2x+1)=3

OpenStudy (solomonzelman):

is it\[Log_2(2x+1)=3~~?\]

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\[\log_ab=c~~~~same~~as~~a^c=b\]

OpenStudy (solomonzelman):

\[2^3=2x+1,~~~~~~8=2x+1,~~~~~~solve~~for~~x.\]

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

Tell me what you get, 3.5 right?

OpenStudy (anonymous):

no not yet

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you much :)

OpenStudy (solomonzelman):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{cyan}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{cyan}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{cyan}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~have~~~a~~~good~~~Thanksgiving!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (anonymous):

Same to you. You are awesome!

OpenStudy (solomonzelman):

Me, ty!